# End-to-end Learnable Clustering for Intent Learning in Recommendation

### 11 Jan 2024 v1

@@ Abstract.
Mining users’ intents plays a crucial role in sequential recommendation. The recent approach, ICLRec, was introduced to extract underlying users’ intents using contrastive learning and clustering. While it has shown effectiveness, the existing method suffers from complex and cumbersome alternating optimization, leading to two main issues. Firstly, the separation of representation learning and clustering optimization within a generalized expectation maximization (EM) framework often results in sub-optimal performance. Secondly, performing clustering on the entire dataset hampers scalability for large-scale industry data. To address these challenges, we propose a novel intent learning method called ELCRec, which integrates representation learning into an End-to-end Learnable Clustering framework for Recommendation. Specifically, we encode users’ behavior sequences and initialize the cluster centers as learnable network parameters. Additionally, we design a clustering loss that guides the networks to differentiate between different cluster centers and pull similar samples towards their respective cluster centers. This allows simultaneous optimization of recommendation and clustering using mini-batch data. Moreover, we leverage the learned cluster centers as self-supervision signals for representation learning, resulting in further enhancement of recommendation performance. Extensive experiments conducted on open benchmarks and industry data validate the superiority, effectiveness, and efficiency of our proposed ELCRec method. Code is available at: https://github.com/yueliu1999/ELCRec.
@@ 1. Introduction
To solve these problems, a novel intention learning method called ELCRec is proposed, as shown in Figure 3. The basic idea of this method is to better capture user intent by unifying sequential representation learning in a End-to-end Llearnable Clustering framework, thereby achieving better underlining recommendedrecommended. Specifically, the user’s behavioral process is first embedded into the latent space. Based on the behavioral embeddings, the cluster centers are then initialized as learnable neural network parameters. Furthermore, we design a clustering loss to guide the network to replace different cluster centers and merge homogeneous samples to cluster centers. This loss function allows our method to optimize the cluster distribution on small batches of data, improves the scalability of our method and solves problem (2). Furthermore, to further improve recommendation performance, the learned cluster centers are assigned as self-monitoring signals for sequential representation learning. The optimizer minimizes both clustering loss and recommendation loss simultaneously. These settings unify representation learning and cluster optimization in an end-to-end framework, improving recommendation performance and simplifying deployment. Therefore, problem (1) is also solved.
@@
  • A intent learning method termed ELCRec is proposed to better mine users’ underlying intents via unifying representation learning and clustering optimization.
  • An end-to-end learnable clustering module is presented to optimize clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment.
  • Extensive experiments demonstrate the capability of ELCRec from six perspectives, including superiority, effectiveness, efficiency, sensitivity, convergence, and visualization.
@@ 2.1. Sequential Recommendation
In recent studies, it has been argued by researchers that users’ intents play a pivotal role in recommendation systems. For instance, Chen et al. (ICLRec, ) introduce a general framework called ICLRec, which utilizes $k$-Means clustering to extract users’ underlying intents. However, the alternating optimization process in ICLRec can be intricate and burdensome, thus hindering scalability and reproducibility. To address this limitation, this paper presents an end-to-end learnable clustering framework for intent learning in recommendation systems.
@@ 2.3. Clustering Algorithm
Figure 3. Overview of our proposed ELCRec. At first, the encoder embeds the users’ historical behavior sequences into the latent space. Based on the embeddings, we initialize the cluster centers as the learnable neural parameters (tensor with gradient). In addition, the clustering distribution can be optimized by the proposed clustering loss with the mini-batch data, thus improving the scalability. Besides, the cluster-assisted contrastive learning is presented to assist the sequential representation learning. By these settings, the clustering optimization and representation learning are unified into an end-to-end framework, improving the recommendation performance and convenience of deployment.
@@ 3. Methodology
In this section, we will present our proposed framework, ELCRec. Firstly, we provide the necessary notations and define the task. Next, we analyze and identify the limitations in existing intent learning methods. Finally, we propose solutions to address these challenges.
@@ 3.3. Problem Analyses
To alleviate these issues, we propose an end-to-end learnable clustering framework for intent learning in recommendation.
@@ 3.4. Proposed Method
To solve the previous problems, we present a new intent learning method termed ELCRec by unifying sequence representation learning into an End-to-end Learnable Clustering framework, for Recommendation. The overall framework of ELCRec is demonstrated in Figure 3. It mainly contains three parts, including sequence encoding, end-to-end learnable cluster module, and cluster-assisted contrastive learning. Next, we introduce them in detail.
@@ 3.4.1. Encoding
3.4.2. End-to-end Learnable Cluster Module
In this section, an End-to-end Learnable Cluster Module (ELCM) is proposed to mine the users’ underlying intents. Concretely, the proposed clustering algorithm groups the users’ behavior embeddings into various clusters, which can represent the latent intents of users.
At first, based on $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$, the cluster centers $\mathbf{C}\in\mathbb{R}^{k\times d}$ are initialized as the learnable neural parameters. Namely, they are the tensors with gradients. Then, we present a clustering loss to train the networks and cluster centers as formulated as follows.
@@ 3.4.2. End-to-end Learnable Cluster Module
Differently, in ICLRec (ICLRec, ), the $k$-Means clustering need to be performed on users’ behavior embeddings $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$. In each E-step, it leads to $\mathcal{O}(t|\mathcal{U}|kd)$ time complexity and $\mathcal{O}(|\mathcal{U}|k+|\mathcal{U}|d+kd)$ space complexity, where $t$ denote the iteration steps of $k$-Means. Both the time complexity and space complexity are linear to the number of users, easily leading to the out-of-memory or the long running time problems in the large-scale industrial data. Our proposed end-to-end learnable cluster module can solve these issues. By summarising previous analyses, the overall time and space complexity of calculating the clustering loss $\mathcal{L}_{\text{cluster}}$ are $\mathcal{O}(bkd+k^{2}d+bd)$ and $\mathcal{O}(bk+bd+kd)$, respectively. They are both linear to the batch size $b$. Therefore, it is scalable and plug-and-play, improving the convenience of deployment. For the evidences, please refer to Section 5. Besides, the learnable cluster centers can unify the representation learning and clustering optimization into an end-to-end framework, achieving better recommendation performance. Experimental evidences are listed in Section 4.2 and Section 4.3. In addition, benefiting from the end-to-end framework, the time and space efficiency are both improved, as shown in Section 4.4.
@@ 3.4.3. Cluster-assisted Contrastive Learning
Additionally, to further collaborate intent learning and sequential representation learning, we conduct the contrastive learning between the user’s behaviors and the learnable intent centers. The intent contrastive loss is formulated as follows.
@@
In this section, the learn clusters are guided to assist the contrastive learning for sequential representation learning. In summary, the objective of cluster-assisted contrastive learning is formulated as $\mathcal{L}_{\text{ccl}}=\mathcal{L}_{\text{seq\_cl}}+\mathcal{L}_{\text{intent\_cl}}$. In this manner, the clustering optimization and representation learning are unified into an end-to-end framework, improving the recommendation performance and convenience of deployment. The effectiveness of CCL is verified in Section 4.3.
@@ 3.4.4. Overall Objective
The neural networks and learnable clusters are trained with multiple tasks, including clustering, cluster-assisted contrastive learning, and next-item prediction. The clustering task aims to capture the users’ underlying intents. Besides, the cluster-assisted contrastive learning aims to collaborate the intent learning and sequential representation learning. In addition, the next-item prediction task is a widely used task for recommendation system (SASRec, ; S3_rec, ; ComiRec, ; ICLRec, ). In summary, the overall objective of the proposed ELCRec is formulated as follows.
@@ 4. Experiment
This section aims to comprehensively evaluate our proposed ELCRec by answering the following research questions.
@@
    Superiority: does ELCRec outperform the existing state-of-the-art sequential recommendation methods?
@@
    Effectiveness: are the proposed end-to-end learnable cluster module and cluster-assisted contrastive learning effective?
@@
    Efficiency: how about the time and memory efficiency of the proposed ELCRec?
@@ 4.1.5. Implementation
For the baselines, we utilized the implementation provided in (ICLRec, ). Our proposed method, ELCRec, was implemented using the PyTorch deep learning platform. In the Transformer encoder, we employed self-attention blocks with 2 attention heads. The latent dimension, denoted as $d$, was set to 64, and the maximum sequence length, denoted as $T$, was set to 50. We utilized the Adam optimizer with a learning rate of 1e-03. The decay rate for the first moment estimate was set to 0.9, and the decay rate for the second moment estimate was set to 0.999. The cluster number, denoted as $k$, was set to 256 for the Yelp and Beauty datasets, and 512 for the Sports and Toys datasets. The trade-off hyper-parameter, denoted as $\alpha$, was set to 1 for the Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. During training, we monitored the training process using the Weights & Biases platform.
@@ 4.2. Superiority
In this section, we aim to answer research question (1) and demonstrate the superiority of the proposed ELCRec method. To be specific, we compare ELCRec with nine state-of-the-art recommendation baselines (BPR_MF, ; GRU4Rec, ; Caser, ; SASRec, ; DSSRec, ; BERT4Rec, ; S3_rec, ; CL4Rec, ; ICLRec, ). Experimental results are the mean values of three runs. As shown in Table 1, the bold and underlined values denote the best and runner-up results, respectively. From these results, we have three conclusions as follows.
@@
    More recently, the intent learning methods (DSSRec, ; ICLRec, ) have been proposed to mine users’ underlying intent to assist recommendation. Motivated by their success, we propose an end-to-end learnable cluster framework named ELCRec for intent learning. Befitting from the strong intent mining capability of ELCRec, it surpasses other intent learning methods. To further verify the superiority of ELCRec, we conduct the $t$-test between the best and runner-up methods. As shown in Figure 1, the most $p$-value is less than 0.05 except HR@5 on the Toys dataset. This indicates our proposed ELCRec significantly outperforms the runner-up method.
@@
Dataset Metric BPR-MF GRU4Rec Caser SASRec DSSRec BERT4Rec S3-Rec CL4SRec ICLRec ELCRec $p$-value
@@ 4.3. Effectiveness
This section is dedicated to addressing research question (2) and evaluating the effectiveness of the End-to-end Learnable Cluster Module (ELCM) and Cluster-assisted Contrastive Learning (CCL). To achieve this, we conducted meticulous ablation studies on four publicly available datasets. Figure 4 illustrates the results, where sub-figures (a), (b), (c), and (d) represent the Sports, Beauty, Toys, and Yelp datasets, respectively. The results are aggregated from four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. In each sub-figure, the labels “B,” “B+CCL,” “B+ELCM,” and “ELCRec” correspond to the baseline, baseline with CCL, baseline with ELCM, and baseline with both CCL and ELCM, respectively. Through the ablation studies, we draw three key conclusions.
@@
Figure 4. Ablation studies of the proposed end-to-end learnable cluster module and the cluster-assisted contrastive learning. The results are the sum of four metrics, including HR@5, HR@20, NDCG@5, and NDCG@20.
@@
    Secondly, we observe that “B+ELCM” can surpass the baseline “B” significantly on all benchmarks. This demonstrates that our proposed end-to-end learnable cluster module helps the model better capture the users’ underlying intents, therefore achieving promising performance.
@@
    Thirdly, the baseline with CCL and ELCM (i.e., “ELCRec”) achieves the best performance on three out of four datasets. It verifies the effectiveness of the combination of these two modules. On the Toys dataset, “ELCRec” can outperform the baseline but achieve worse performance than “B+CCL” and “B+ELCM”. This phenomenon indicates it is worth doing research on the better collaboration of these two modules.
To summarize, the extensive ablation studies verify the effectiveness of modules in the proposed ELCRec.
@@ 4.4. Efficiency
In this section, we test the efficiency of ELCRec and answer the research question (3). Specifically, the efficiency contains two perspectives, including training time costs and GPU memory costs. We summarize two observations as follows.
@@
    The results of training time costs are listed in Table 2. We find that our proposed ELCRec can speed up ICLRec on three out of four datasets. Overall, on four datasets, the training time is decreased by 7.18% on average.
@@
    The results of GPU memory costs are listed in Table 3. The results demonstrate that the GPU memory costs of our ELCRec are lower than that of ICLRec on four datasets. On average, the GPU memory costs are decreased by 9.58%.
In summary, both the time efficiency and memory efficiency of our proposed ELCRec are verified on four benchmarks.
@@ 5. Application
Benefiting from the versatility of the proposed ELCRec, we aim to apply it to the real-world industrial scenario. The applied scenario is the live-streaming recommendation on the front page of the Alipay app. The UV and PV of this application is about 50 million and 130 million, respectively. Most users are new to this application, therefore leading to the sparsity of users’ behaviors. To solve this cold-start problem in the recommendation system, we adopt the proposed end-to-end learnable cluster module to group users and recommend. Firstly, due to the sparsity of users’ behaviors, we replace the users’ behavior with the users’ activities and model them via the multi-gate mixture-of-expert (MMOE) (MMOE, ). Then, the end-to-end learnable cluster module is adopted to group the users into various clusters. Through this module, the high-activity users and new users are grouped into different clusters, alleviating the cold-start issue and assisting in better recommendations. Eventually, the networks are trained with the multi-task training strategy. Next, we introduce the results and case studies on the industrial data.

### 24 Jan 2024 v2

# Online Differentiable Clustering for Intent Learning in Recommendation

@@ Abstract.
Mining users’ intents plays a crucial role in sequential recommendation. The recent approach, ICLRec, was introduced to extract underlying users’ intents using contrastive learning and clustering. While it has shown effectiveness, the existing method suffers from complex and cumbersome alternating optimization, leading to two main issues. Firstly, the separation of representation learning and clustering optimization within a generalized expectation maximization (EM) framework often results in sub-optimal performance. Secondly, performing clustering on the entire dataset hampers scalability for large-scale industry data. To address these challenges, we propose a novel intent learning method called ODCRec, which integrates representation learning into an Online Differentiable Clustering framework for Recommendation. Specifically, we encode users’ behavior sequences and initialize the cluster centers as differentiable network parameters. Additionally, we design a clustering loss that guides the networks to differentiate between different cluster centers and pull similar samples towards their respective cluster centers. This allows simultaneous optimization of recommendation and clustering using mini-batch data. Moreover, we leverage the learned cluster centers as self-supervision signals for representation learning, resulting in further enhancement of recommendation performance. Extensive experiments conducted on open benchmarks and industry data validate the superiority, effectiveness, and efficiency of our proposed ODCRec method. Code is available at: https://github.com/yueliu1999/ELCRec.
@@ 1. Introduction
To solve these problems, a novel intention learning method called ODCRec is proposed, as shown in Figure 3. The basic idea of this method is to better capture user intent by unifying sequential representation learning in a Online Differentiable Clustering framework, thereby achieving better underlining recommendedrecommended. Specifically, the user’s behavioral process is first embedded into the latent space. Based on the behavioral embeddings, the cluster centers are then initialized as differentiable neural network parameters. Furthermore, we design a clustering loss to guide the network to replace different cluster centers and merge homogeneous samples to cluster centers. This loss function allows our method to optimize the cluster distribution on small batches of data, improves the scalability of our method and solves problem (2). Furthermore, to further improve recommendation performance, the learned cluster centers are assigned as self-monitoring signals for sequential representation learning. The optimizer minimizes both clustering loss and recommendation loss simultaneously. These settings unify representation learning and cluster optimization in a unified framework, improving recommendation performance and simplifying deployment. Therefore, problem (1) is also solved.
@@
  • A intent learning method termed ODCRec is proposed to better mine users’ underlying intents via unifying representation learning and clustering optimization.
  • An online differentiable clustering module is presented to optimize clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment.
  • Extensive experiments demonstrate the capability of ODCRec from six perspectives, including superiority, effectiveness, efficiency, sensitivity, convergence, and visualization.
@@ 2.1. Sequential Recommendation
In recent studies, it has been argued by researchers that users’ intents play a pivotal role in recommendation systems. For instance, Chen et al. (ICLRec, ) introduce a general framework called ICLRec, which utilizes $k$-Means clustering to extract users’ underlying intents. However, the alternating optimization process in ICLRec can be intricate and burdensome, thus hindering scalability and reproducibility. To address this limitation, this paper presents an online differentiable clustering framework for intent learning in recommendation systems.
@@ 2.3. Clustering Algorithm
Figure 3. Overview of our proposed ODCRec. At first, the encoder embeds the users’ historical behavior sequences into the latent space. Based on the embeddings, we initialize the cluster centers as the differentiable neural parameters (tensor with gradient). In addition, the clustering distribution can be optimized by the proposed clustering loss with the mini-batch data, thus improving the scalability. Besides, the cluster-assisted contrastive learning is presented to assist the sequential representation learning. By these settings, the clustering optimization and representation learning are unified into a unified framework, improving the recommendation performance and convenience of deployment.
@@ 3. Methodology
In this section, we will present our proposed framework, ODCRec. Firstly, we provide the necessary notations and define the task. Next, we analyze and identify the limitations in existing intent learning methods. Finally, we propose solutions to address these challenges.
@@ 3.3. Problem Analyses
To alleviate these issues, we propose an online differentiable clustering framework for intent learning in recommendation.
@@ 3.4. Proposed Method
To solve the previous problems, we present a new intent learning method termed ODCRec by unifying sequence representation learning into an Online Differentiable Clustering framework, for Recommendation. The overall framework of ODCRec is demonstrated in Figure 3. It mainly contains three parts, including sequence encoding, online differentiable cluster module, and cluster-assisted contrastive learning. Next, we introduce them in detail.
@@ 3.4.1. Encoding
3.4.2. Online Differentiable Cluster Module
In this section, an Online Differentiable Cluster Module (ELCM) is proposed to mine the users’ underlying intents. Concretely, the proposed clustering algorithm groups the users’ behavior embeddings into various clusters, which can represent the latent intents of users.
At first, based on $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$, the cluster centers $\mathbf{C}\in\mathbb{R}^{k\times d}$ are initialized as the differentiable neural parameters. Namely, they are the tensors with gradients. Then, we present a clustering loss to train the networks and cluster centers as formulated as follows.
@@ 3.4.2. Online Differentiable Cluster Module
Differently, in ICLRec (ICLRec, ), the $k$-Means clustering need to be performed on users’ behavior embeddings $\mathbf{H}\in\mathbb{R}^{|\mathcal{U}|\times d}$. In each E-step, it leads to $\mathcal{O}(t|\mathcal{U}|kd)$ time complexity and $\mathcal{O}(|\mathcal{U}|k+|\mathcal{U}|d+kd)$ space complexity, where $t$ denote the iteration steps of $k$-Means. Both the time complexity and space complexity are linear to the number of users, easily leading to the out-of-memory or the long running time problems in the large-scale industrial data. Our proposed online differentiable cluster module can solve these issues. By summarising previous analyses, the overall time and space complexity of calculating the clustering loss $\mathcal{L}_{\text{cluster}}$ are $\mathcal{O}(bkd+k^{2}d+bd)$ and $\mathcal{O}(bk+bd+kd)$, respectively. They are both linear to the batch size $b$. Therefore, it is scalable and plug-and-play, improving the convenience of deployment. For the evidences, please refer to Section 5. Besides, the differentiable cluster centers can unify the representation learning and clustering optimization into a unified framework, achieving better recommendation performance. Experimental evidences are listed in Section 4.2 and Section 4.3. In addition, benefiting from the unified framework, the time and space efficiency are both improved, as shown in Section 4.4.
@@ 3.4.3. Cluster-assisted Contrastive Learning
Additionally, to further collaborate intent learning and sequential representation learning, we conduct the contrastive learning between the user’s behaviors and the differentiable intent centers. The intent contrastive loss is formulated as follows.
@@
In this section, the learn clusters are guided to assist the contrastive learning for sequential representation learning. In summary, the objective of cluster-assisted contrastive learning is formulated as $\mathcal{L}_{\text{ccl}}=\mathcal{L}_{\text{seq\_cl}}+\mathcal{L}_{\text{intent\_cl}}$. In this manner, the clustering optimization and representation learning are unified into a unified framework, improving the recommendation performance and convenience of deployment. The effectiveness of CCL is verified in Section 4.3.
@@ 3.4.4. Overall Objective
The neural networks and differentiable clusters are trained with multiple tasks, including clustering, cluster-assisted contrastive learning, and next-item prediction. The clustering task aims to capture the users’ underlying intents. Besides, the cluster-assisted contrastive learning aims to collaborate the intent learning and sequential representation learning. In addition, the next-item prediction task is a widely used task for recommendation system (SASRec, ; S3_rec, ; ComiRec, ; ICLRec, ). In summary, the overall objective of the proposed ODCRec is formulated as follows.
@@ 4. Experiment
This section aims to comprehensively evaluate our proposed ODCRec by answering the following research questions.
@@
    Superiority: does ODCRec outperform the existing state-of-the-art sequential recommendation methods?
@@
    Effectiveness: are the proposed online differentiable cluster module and cluster-assisted contrastive learning effective?
@@
    Efficiency: how about the time and memory efficiency of the proposed ODCRec?
@@ 4.1.5. Implementation
For the baselines, we utilized the implementation provided in (ICLRec, ). Our proposed method, ODCRec, was implemented using the PyTorch deep learning platform. In the Transformer encoder, we employed self-attention blocks with 2 attention heads. The latent dimension, denoted as $d$, was set to 64, and the maximum sequence length, denoted as $T$, was set to 50. We utilized the Adam optimizer with a learning rate of 1e-03. The decay rate for the first moment estimate was set to 0.9, and the decay rate for the second moment estimate was set to 0.999. The cluster number, denoted as $k$, was set to 256 for the Yelp and Beauty datasets, and 512 for the Sports and Toys datasets. The trade-off hyper-parameter, denoted as $\alpha$, was set to 1 for the Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. During training, we monitored the training process using the Weights & Biases platform.
@@ 4.2. Superiority
In this section, we aim to answer research question (1) and demonstrate the superiority of the proposed ODCRec method. To be specific, we compare ODCRec with nine state-of-the-art recommendation baselines (BPR_MF, ; GRU4Rec, ; Caser, ; SASRec, ; DSSRec, ; BERT4Rec, ; S3_rec, ; CL4Rec, ; ICLRec, ). Experimental results are the mean values of three runs. As shown in Table 1, the bold and underlined values denote the best and runner-up results, respectively. From these results, we have three conclusions as follows.
@@
    More recently, the intent learning methods (DSSRec, ; ICLRec, ) have been proposed to mine users’ underlying intent to assist recommendation. Motivated by their success, we propose an online differentiable cluster framework named ODCRec for intent learning. Befitting from the strong intent mining capability of ODCRec, it surpasses other intent learning methods. To further verify the superiority of ODCRec, we conduct the $t$-test between the best and runner-up methods. As shown in Figure 1, the most $p$-value is less than 0.05 except HR@5 on the Toys dataset. This indicates our proposed ODCRec significantly outperforms the runner-up method.
@@
Dataset Metric BPR-MF GRU4Rec Caser SASRec DSSRec BERT4Rec S3-Rec CL4SRec ICLRec ODCRec $p$-value
@@ 4.3. Effectiveness
This section is dedicated to addressing research question (2) and evaluating the effectiveness of the Online Differentiable Cluster Module (ELCM) and Cluster-assisted Contrastive Learning (CCL). To achieve this, we conducted meticulous ablation studies on four publicly available datasets. Figure 4 illustrates the results, where sub-figures (a), (b), (c), and (d) represent the Sports, Beauty, Toys, and Yelp datasets, respectively. The results are aggregated from four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. In each sub-figure, the labels “B,” “B+CCL,” “B+ELCM,” and “ODCRec” correspond to the baseline, baseline with CCL, baseline with ELCM, and baseline with both CCL and ELCM, respectively. Through the ablation studies, we draw three key conclusions.
@@
Figure 4. Ablation studies of the proposed online differentiable cluster module and the cluster-assisted contrastive learning. The results are the sum of four metrics, including HR@5, HR@20, NDCG@5, and NDCG@20.
@@
    Secondly, we observe that “B+ELCM” can surpass the baseline “B” significantly on all benchmarks. This demonstrates that our proposed online differentiable cluster module helps the model better capture the users’ underlying intents, therefore achieving promising performance.
@@
    Thirdly, the baseline with CCL and ELCM (i.e., “ODCRec”) achieves the best performance on three out of four datasets. It verifies the effectiveness of the combination of these two modules. On the Toys dataset, “ODCRec” can outperform the baseline but achieve worse performance than “B+CCL” and “B+ELCM”. This phenomenon indicates it is worth doing research on the better collaboration of these two modules.
To summarize, the extensive ablation studies verify the effectiveness of modules in the proposed ODCRec.
@@ 4.4. Efficiency
In this section, we test the efficiency of ODCRec and answer the research question (3). Specifically, the efficiency contains two perspectives, including training time costs and GPU memory costs. We summarize two observations as follows.
@@
    The results of training time costs are listed in Table 2. We find that our proposed ODCRec can speed up ICLRec on three out of four datasets. Overall, on four datasets, the training time is decreased by 7.18% on average.
@@
    The results of GPU memory costs are listed in Table 3. The results demonstrate that the GPU memory costs of our ODCRec are lower than that of ICLRec on four datasets. On average, the GPU memory costs are decreased by 9.58%.
In summary, both the time efficiency and memory efficiency of our proposed ODCRec are verified on four benchmarks.
@@ 5. Application
Benefiting from the versatility of the proposed ODCRec, we aim to apply it to the real-world industrial scenario. The applied scenario is the live-streaming recommendation on the front page of the Alipay app. The UV and PV of this application is about 50 million and 130 million, respectively. Most users are new to this application, therefore leading to the sparsity of users’ behaviors. To solve this cold-start problem in the recommendation system, we adopt the proposed online differentiable cluster module to group users and recommend. Firstly, due to the sparsity of users’ behaviors, we replace the users’ behavior with the users’ activities and model them via the multi-gate mixture-of-expert (MMOE) (MMOE, ). Then, the online differentiable cluster module is adopted to group the users into various clusters. Through this module, the high-activity users and new users are grouped into different clusters, alleviating the cold-start issue and assisting in better recommendations. Eventually, the networks are trained with the multi-task training strategy. Next, we introduce the results and case studies on the industrial data.